# Learning to Predict Program Execution by Modeling Dynamic Dependency on Code Graphs

### v2

@@ Abstract
Predicting program behavior without execution is an essential and challenging task in software engineering. Traditional models often struggle to capture dynamic dependencies and interactions within code. This paper introduces a novel machine learning-based framework called CodeFlow, which predicts code coverage and detects runtime errors through Dynamic Dependencies Learning. Utilizing control flow graphs (CFGs), CodeFlow represents all possible execution paths and the relationships between different statements, offering a comprehensive understanding of program behavior. It constructs CFGs to depict execution paths and learns vector representations for CFG nodes, capturing static control-flow dependencies. Additionally, it learns dynamic dependencies through execution traces, which reflect the impacts among statements during execution. This approach enables accurate prediction of code coverage and identification of runtime errors. Empirical evaluations show significant improvements in code coverage prediction accuracy and effective localization of runtime errors, surpassing current models.
@@ I Introduction
To show CodeFlow’s usefulness, we leverage it to build a tool to statically detect runtime errors in (in)complete code snippets. Platforms like Stack Overflow (S/O) are invaluable resources for developers facing technical issues. However, S/O code snippets may contain hidden defects, runtime errors, and security vulnerabilities, posing potential risks to applications that integrate them. Integrating a code snippet into a codebase for analysis is time-consuming and risky. To mitigate these risks, it is essential to directly analyze online code snippets to reason about their behaviors. However, the key challenge is the incompleteness nature of online code snippets. To do so, we use an LLM to act as a fuzzer, generating inputs to detect runtime errors in a given code snippet. Each input is used in the code snippet, which is then fed into CodeFlow to predict the code coverage. If the code coverage stops unexpectedly and never reaches an exit point, CodeFlow will locate the error.
We conducted an empirical evaluation on CodeFlow. Our findings indicate that it significantly improves code coverage prediction, runtime error detection, and bug localization compared to existing models. Specifically, CodeFlow achieves an accuracy of 75.24% in matching code coverage exactly, outperforming GPT-4o at 68.13%. For branch coverage, CodeFlow reaches 87.88%, significantly higher than GPT-4o’s 78.75%. In runtime error detection, CodeFlow attains a high accuracy of 97.51%, far exceeding the performance of other models. Additionally, CodeFlow maintains high accuracy even on incomplete code snippets, demonstrating its generalization capability. Furthermore, CodeFlow proves highly effective in supporting fuzz testing, particularly in scenarios involving incomplete code snippets where traditional execution is not feasible.
@@ II-A Example and Observations
Conversely, CodeExecutor [9] correctly skips the else branch by understanding the intermediate values of x throughout the execution trace. However, it performs poorly with complex code requiring multiple iterations to update variable values. Error propagation frequently occurs, leading to incorrect code coverage predictions.
@@ II-A3 Observation 3: Information Loss in Repeated Loops
For loops that run numerous iterations, predicting the outcome based on line-by-line variable states often leads to incorrect results. For example, the for loop on line 11 runs 100 iterations, updating x each time (lines 12-13). While the LLM correctly predicts that the loop will execute all iterations, it fails to understand the cumulative effect on x. This is because the LLM only processes static information, whereas the for loop requires information from the last line in the body to be fed back into the loop. After the loop, the value of x is used to decide which branch of the if-elif-else statement (lines 14-19) will be executed. The LLM incorrectly predicts the outcome because it does not account for the aggregated change in x. Consequently, it fails to accurately simulate the dynamic changes in variable states over multiple iterations.
Figure 2: Control Flow Graph for code in Fig. 1
@@ VI Dynamic Dependencies Learning
With function $f$ in (3) is aggregation function from previous states to summarize history information. For usual language modeling task, the function $f$ normally takes form of $f(h_{0},h_{1},\cdots,h_{i-1})=h_{i-1}$.
But considering the following characteristics from our problem, we propose an adaptation to the original LSTM which was originally used for language modeling $P(t_{i}|t_{0},t_{1},\cdots,t_{i-1})$
@@
2. Forward and backward passing: in our CFG, a loop is broken down into condition node, body, and step node (e.g: $\text{counter}+=1$). The forward edges are just ordinary increasing order, $n_{i}\to n_{j}$ but what is special is an additional backward edge from the loop step node to the condition node. To propagate the information from the step node $n_{j}$ to condition node $n_{i}$ with $i<j$ but not to mix up the recurrent relations of execution. We update the state $h_{i}$ by $h_{j}$ for this special backward edge only by information of node embedding $\overline{x_{j}}$. By updating only the information but not the state at node $n_{j}$ which is not yet computed by forward order, we can combine both forward and backward propagation by only updating $h_{i}=LSTM(h_{j})$ with $i<j$ instead of bidirectional like $BiLSTM$.
3. Binary soft decision branching: we have processed a condition (including loop condition) node $n_{i}$ will only connect to two nodes $n_{j}$ and $n_{k}$ with $i<j,k$ by two forward edges. And each of these edges is the only incoming edge to $n_{j}$ and $n_{k}$. The original computation of node $n_{j}$’s hidden state is $h_{j}^{*}=A_{forward}H=[h_{0},\cdots,h_{i-1},\mathbb{0},\cdots,\mathbb{0}]=h_{i}$, similarly, $h_{j}^{*}=h_{i}$. With that, we allow information to pass to both possible branches while they are complemented in real code execution. Thus in order to model the branching behavior in the condition node, we will charge $A$ dynamically based on the current hidden state $h_{i}$. $A[i,j]=0$ if $Average(h_{i})\geq 0$ and $A[i,k]=0$ if $Average(h_{i})<0$ with assumption $j<k$. This will force the weights to adaptively produce reasonable $h_{i}$ to make a correct branch decision. A similar approach is also taken by [15], but the major difference is that they add additional parameter complexity to learn this soft dynamic branching while we focus more on the efficiency by setting branching conditions depending only on the current hidden state.
@@ VIII Empirical Evaluation
RQ1. [Code Coverage Prediction Accuracy]: How well does CodeFlow predict code coverage for (in)complete code in Python?
@@ IX-A1 Datasets
However, CodeNetMut lacks a sufficient number of Python files with extensive conditional statements, which are crucial for training the model on the behavior of conditional branching. To address that, we supplemented CodeNetMut with an additional dataset generated using Gemini-API. This synthetic dataset comprises approximately 11,668 Python code snippets, each featuring diverse and complex types of statements. For a snippet, we generate the CFG in which we track the nodes, modeling individual statements or blocks of code, forward edges (representing the normal control flow from a statement to another), and backward edges (representing the control flow from the last line of a loop back to the loop node.
To build the ground truth in training, we use the trace library from Python to record which lines of code are executed when running each snippet and record the final code coverage.
@@ IX-A3 Evaluation Metrics
To evaluate the models, we use the following key metrics.
@@
+ Precision (P), Recall (R), F1-Score (F1): Precision is the ratio of correctly predicted executed nodes (true positives) to the total number of nodes predicted as executed (true positives + false positives). Recall is the ratio of correctly predicted executed nodes (true positives) to the total number of actually executed nodes (true positives + false negatives). The F1-Score is the harmonic mean of Precision and Recall.
@@ IX-B Empirical Results
TABLE I: Code Coverage Prediction Comparison (RQ1).
Model EM (%) BC (%) P R F1
CodeEx 18.83 31.34 0.94 0.47 0.63
CFGNN 45.53 76.56 0.92 0.91 0.92
LLMs
   Gemini 56.17 74.96 0.87 0.97 0.92
   Claude 64.94 77.30 0.96 0.94 0.95
   GPT-4 68.13 78.75 0.96 0.96 0.96
CodeFlow 75.24 87.88 0.97 0.97 0.97
As seen in Table I, CodeFlow outperforms existing models across all key metrics.
@@ IX-B1 Exact-Matching
CodeFlow achieves an exact matching accuracy of 75.24%, which is higher than the best-performing LLM, GPT-4o, at 68.13%. Notably, CodeFlow accomplishes this with far fewer parameters (1.3 million in total) compared to LLMs like GPT-4o (over 1 trillion) and Claude (175 billion), which suggests a more efficient architecture for this task and underscores its practicality for the scenarios where computational resources may be limited.
@@
IX-B2 Branch-coverage Matching
CodeFlow achieves 87.88% correctness, outperforming the next best baseline, GPT-4o, by nearly 10%. This margin highlights the efficacy of using CFGs to capture the complex behaviors of loops and conditional branches. With CFG modeling the intricate decision points within code, CodeFlow understands the cumulative effect of variable changes over multiple iterations, leading to more accurate branch predictions. In contrast, models like CodeExecutor fail in this aspect because they do not adequately handle dynamic execution changes, leading to predictions that do not align well with actual execution paths involving loops and conditions.
CodeFlow also excels in precision, recall, and F1-score metrics, achieving 97% in all three categories. These values are slightly higher than those of the best LLMs (Claude and GPT-4o), which both scored 96% in precision and recall. The high precision indicates that CodeFlow is very effective in identifying executed lines without falsely predicting non-executed lines as executed (minimizing false positives). The recall score reflects the model’s ability to capture all executed lines without missing any (minimizing false negatives). Regarding the LLMs, note that half of the test dataset comprises solutions from the CodeNet Project published online since 2022, potentially giving them an advantage if they were trained on this data. Despite this, CodeFlow still outperforms them.
IX-B3 Continuity in Predicting Code Coverage
Continuity means that a model should predict nodes or lines of code that are close together or logically lead to each other, rather than randomly jumping to distant nodes while ignoring intermediate nodes in between. This is a common issue observed with existing methods. CodeExecutor relies heavily on exact execution traces during training, which causes it to miss the dependencies between lines of code. As seen in Fig. 6a, CodeExecutor skips one node and jumps directly to the next node, resulting in a discontinuous path in the CFG.
@@ IX-B3 Continuity in Predicting Code Coverage
LLMs, such as GPT-4o, also face challenges in maintaining continuity. They sometimes fail to understand the dynamic relations between lines and may skip essential steps in sequences, e.g., the elif in an if-elif-else structure. This is depicted in Fig. 6c, where the model predicts nodes without capturing the logical sequence. This occurs because LLMs rely primarily on next-token probability, predicting the most likely subsequent token based on the preceding context. This approach can result in misunderstandings of the code structure, especially concerning dynamic behaviors and state changes across iterations. Thus, while having high precision, recall, and F1-score, they have lower exact-matching accuracies.
@@
IX-B4 Incomplete Code
To our knowledge, there is no existing runtime error dataset of incomplete Python code snippets with the respective complete versions (to build the ground truth). Thus, to test its capability with incomplete code, specifically when built-in and external libraries are not imported, we trained and tested CodeFlow on a dataset where we removed all import statements of built-in/external libraries or method/class declarations. This ensures the code’s semantic integrity remains intact including keeping all function/API calls within the code. Despite the absence of explicit import statements, CodeFlow achieved similar results as in Table I (not shown) because it learned the semantic meaning of these libraries during training, allowing it to understand and predict code behavior based on the nodes that use the library functions. Similar performance was observed in LLMs (GPT-4o, Claude, and Gemini), as their extensive pre-training includes library semantics. CodeExecutor, which tracks intermediate values in the execution trace, and CFGNN, which models control flow, also performed well without import information.
@@ X-A Data Collection and Evaluation Metrics
For the purpose of evaluating runtime error detection accuracy, we utilized a dataset in addition to the one in RQ1. Specifically, we employed the FixEval dataset, which comprises 2,066 unique problems with a total of 277,262 submissions of Python code snippets. From this dataset, test cases were obtained for 800 problems from the CodeNet dataset [16]. Each of these Python snippets, when executed with its respective input, leads to a runtime error.
After filtering, we obtained 6,437 submissions across the 800 problems, which we added to our dataset. This combined dataset, referred to as the Complete Runtime-Error Dataset, includes both code snippets that terminate normally and those that encounter runtime errors.
To evaluate CodeFlow in detecting runtime errors, in addition to the metrics in RQ1, we use Runtime Error Detection Accuracy (EDA) to measure the accuracy of a model correctly predicting if a code snippet has an error.
@@ X-B Empirical Results
As presented in Table II, Claude achieved the highest precision score of 0.98, indicating its strong ability to correctly identify runtime errors when they are present. High precision means Claude makes very few false positive predictions, thus demonstrating its reliability in accurately pinpointing real runtime errors. However, Claude’s recall score of 0.89, while still respectable, is lower than that of GPT-4o. This suggests that Claude may miss some actual runtime errors, indicating that it is more conservative in its error detection.
In contrast, GPT-4o achieved the highest recall score of 0.99, which highlights its effectiveness in identifying nearly all actual runtime errors. This high recall shows GPT-4o’s thoroughness in detecting errors. However, GPT-4o’s precision score of 0.71 indicates a higher rate of false positives compared to Claude, meaning it sometimes incorrectly flags non-buggy code snippets as erroneous. This suggests that GPT-4o could have more false alarms.
@@
As seen, CodeFlow exhibits a balanced performance with high scores across all metrics: a precision of 0.96, a recall of 0.94, and an F1-score of 0.95. This balance indicates that CodeFlow not only accurately detects a high proportion of actual runtime errors but also minimizes false positives. The overall detection accuracy of 97.51% underscores CodeFlow’s superior capability in statically identifying runtime errors without execution.

### v3

# CodeFlow: Program Behavior Prediction with Dynamic Dependencies Learning

@@ Abstract
Predicting program behavior without execution is a critical task in software engineering. Existing models often fall short in capturing the dynamic dependencies among program elements. To address this, we present CodeFlow, a novel machine learning-based approach that predicts code coverage and detects runtime errors by learning both static and dynamic dependencies within the code. By using control flow graphs (CFGs), CodeFlow effectively represents all possible execution paths and the statistic relations between different statements, providing a more comprehensive understanding of program behaviors. CodeFlow constructs CFGs to represent possible execution paths and learns vector representations (embeddings) for CFG nodes, capturing static control-flow dependencies. Additionally, it learns dynamic dependencies by leveraging execution traces, which reflect the impacts among statements during execution. This combination enables CodeFlow to accurately predict code coverage and identify runtime errors. Our empirical evaluation demonstrates that CodeFlow significantly improves code coverage prediction accuracy and effectively localizes runtime errors, outperforming state-of-the-art models.
Index Terms:
AI4SE, Code Coverage Analysis, Runtime Error Detection, Control Flow Graph
@@ I Introduction
To show CodeFlow’s usefulness in analyzing code coverage and dynamic behaviors of (in)complete code snippets, we use it to build a tool to statically detect runtime errors in both (in)complete snippets. Platforms like Stack Overflow (S/O) are invaluable resources for developers facing technical issues. However, S/O code snippets may contain hidden defects, runtime errors, and security vulnerabilities, posing potential risks to applications that integrate them [12, 13, 14, 15]. It is essential to directly analyze online code snippets to reason on their behaviors. Such vulnerabilities can crash processes or pose security risks if executed without prior analysis. The rationale for early detection is that the execution of such adapted code is unsafe due to the presence of pre-existing vulnerabilities in online code. However, the key challenge is the incompleteness of online code snippets. This incompleteness may also arise from issues such as incompatible libraries or version mismatches (e.g., CUDA incompatibility), which prevent direct execution. Thus, predicting code coverage for incomplete code thus becomes a safer alternative. To do so, we use an LLM to act as a fuzzer, generating inputs to detect runtime errors in a given code snippet. Each input is used in the code snippet, which is then fed into CodeFlow to predict the code coverage. If the code coverage stops unexpectedly and never reaches an exit point, CodeFlow will locate the error.
We conducted an empirical evaluation on CodeFlow. Our findings indicate that it significantly improves code coverage prediction, runtime error detection, and bug localization compared to existing models. Specifically, CodeFlow achieves an accuracy of 75.24% in matching code coverage exactly, outperforming GPT-4o at 68.13%. For branch coverage, CodeFlow reaches 87.88%, significantly higher than GPT-4o’s 78.75%. In runtime error detection, CodeFlow attains a high accuracy of 97.51%, exceeding the performance of other models. Moreover, CodeFlow maintains high accuracy even on incomplete code snippets, demonstrating its generalization capability. CodeFlow also proves highly effective in supporting fuzz testing, particularly in scenarios involving incomplete code snippets where traditional execution is not feasible.
@@ II-A Example and Observations
CodeExecutor [9] correctly skips the else branch by capturing the intermediate values of x throughout the execution trace. However, it performs poorly with complex code requiring multiple iterations to update variable values. Error propagation frequently occurs, leading to incorrect coverage predictions.
Figure 2: Control Flow Graph for code in Fig. 1
@@ II-A3 Observation 3: Information Loss in Repeated Loops
For loops running several iterations, predicting the outcome based on line-by-line variable states often leads to incorrect results. For example, the loop on line 11 runs 100 iterations, updating x each time (lines 12-13). While the LLM correctly predicts that the loop will execute all iterations, it fails to understand the cumulative effect on x. This is because the LLM only processes static information, whereas the for loop requires information from the last line in the body to be fed back into the loop. After the loop, the value of x is used to decide which branch of the if-elif-else statement (lines 14-19) will be executed. The LLM incorrectly predicts the outcome because it does not account for the aggregated change in x. Consequently, it fails to accurately simulate the dynamic changes in variable states over multiple iterations.
@@ VI Dynamic Dependencies Learning
With function $f$ in (3) is aggregation function from previous states to summarize history information. For usual language modeling task, the function $f$ normally takes form of $f(h_{0},h_{1},\cdots,h_{i-1})=h_{i-1}$. But considering the following characteristics from our problem, we propose an adaptation to the original LSTM which was originally used for language modeling $P(t_{i}|t_{0},t_{1},\cdots,t_{i-1})$:
@@
2. Forward and backward passing: in our CFG, a loop is broken down into condition node, body, and step node (e.g: $\text{counter}+=1$). The forward edges are in an increasing order, $n_{i}\to n_{j}$ but what is special is an additional backward edge from the loop step node to the condition one. To propagate the information from the step node $n_{j}$ to condition node $n_{i}$ with $i$$<$$j$ but not to mix up the recurrent relations of execution. We update the state $h_{i}$ by $h_{j}$ for this special backward edge only by information of node embedding $\overline{x_{j}}$. By updating only the information but not the state at node $n_{j}$ which is not yet computed by forward order, we can combine both forward and backward propagation by only updating $h_{i}=LSTM(h_{j})$ with $i$$<$$j$ instead of bidirectional like $BiLSTM$.
3. Binary soft decision branching: we have processed a condition (including loop condition) node $n_{i}$ will only connect to two nodes $n_{j}$ and $n_{k}$ with $i<j,k$ by two forward edges. And each of these edges is the only incoming edge to $n_{j}$ and $n_{k}$. The original computation of node $n_{j}$’s hidden state is $h_{j}^{*}=A_{forward}H=[h_{0},\cdots,h_{i-1},\mathbb{0},\cdots,\mathbb{0}]=h_{i}$, similarly, $h_{j}^{*}=h_{i}$. With that, we allow information to pass to both possible branches while they are complemented in real code execution. Thus in order to model the branching behavior in the condition node, we will charge $A$ dynamically based on the current hidden state $h_{i}$. $A[i,j]=0$ if $Average(h_{i})\geq 0$ and $A[i,k]=0$ if $Average(h_{i})<0$ with assumption $j<k$. This will force the weights to adaptively produce reasonable $h_{i}$ to make a correct branch decision. A similar approach is taken by [19], but the major difference is that they add additional parameter complexity to learn this soft dynamic branching while we focus more on the efficiency by setting branching conditions depending only on the current hidden state.
@@ VIII Empirical Evaluation
RQ1. [Coverage Prediction Accuracy]: How well does CodeFlow predict code coverage for (in)complete code?
@@
TABLE I: Code Coverage Prediction Comparison (RQ1).
Model Complete Code Incomplete Code
EM (%) BC (%) P R F1 EM (%) BC (%) P R F1
CodeExecutor 18.83 31.34 0.94 0.47 0.70 10.45 25.50 0.90 0.42 0.66
CFGNN 45.53 76.56 0.92 0.91 0.92 44.32 77.10 0.90 0.89 0.90
LLMs
   Gemini 56.17 74.96 0.87 0.97 0.92 59.10 73.85 0.88 0.95 0.91
   Claude 64.94 77.30 0.96 0.94 0.95 66.50 79.00 0.95 0.93 0.94
   GPT-4 68.13 78.75 0.96 0.96 0.96 67.75 80.20 0.96 0.95 0.96
CodeFlow 75.24 87.88 0.97 0.97 0.97 76.50 86.95 0.96 0.98 0.97
@@ IX-A1 Datasets
However, CodeNetMut lacks a sufficient number of Python files with extensive conditional statements, which are crucial for training the model on conditional branching. To address this, we supplemented CodeNetMut with an additional dataset generated using Gemini-API. This synthetic dataset comprises approximately 11,668 Python code snippets, each featuring diverse and complex statements. The code snippets vary in size, with the largest containing 146 lines of code and a mean length of 13. Over 4,500 Python code snippets (23%) have a Cyclomatic Complexity above 10, being classified as complex and challenging to test [21]. For each snippet, we generate the CFG, tracking nodes, forward edges (normal control flow), and backward edges (loop control flow). The dataset was split 80:20 into training and testing sets.
To build the ground truth in training, we use the trace library from Python to record the code coverage. In addition, we created the Incomplete Code dataset by removing all import statements and external file references from each snippet.
@@ IX-A3 Evaluation Metrics
We use the following key metrics.
@@
+ Precision (P), Recall (R), F1-Score (F1): These metrics are determined by consolidating all nodes in the test set into a unified dataset and calculating the metrics for that. Precision measures the proportion of nodes predicted as covered that were actually covered during execution, while recall reflects the proportion of executed nodes correctly predicted as covered. F1-Score is the harmonic mean of precision and recall.
@@ IX-B Empirical Results
As seen in Table I, CodeFlow outperforms existing models across all metrics for both complete and incomplete code.
@@ IX-B1 Exact-Matching
For complete code, it achieves an exact matching accuracy of 75.24%, which is higher than the best-performing LLM, GPT-4o, at 68.13%. Notably, CodeFlow accomplishes this with far fewer parameters (1.3 million in total) compared to LLMs like GPT-4o (over 1 trillion) and Claude (175 billion), which suggests a more efficient architecture for this task and underscores its practicality for the scenarios where computational resources may be limited.
IX-B2 Branch-coverage Matching
Our model achieves 87.9% correctness, outperforming the best baseline, GPT-4o, by nearly 10%. This highlights the efficacy of using CFGs to capture the complex behaviors of loops and conditional branches. With CFG modeling the intricate decision points within code, CodeFlow understands the cumulative effect of variable changes over multiple iterations, leading to more accurate branch predictions. In contrast, models like CodeExecutor fail in this aspect because they do not adequately handle dynamic execution changes, leading to predictions that do not align well with actual execution paths involving loops and conditions.
CodeFlow achieves 97% in all three metrics, slightly surpassing the best LLMs (Claude and GPT-4o), which scored 96%. Its high precision shows effectiveness in identifying executed lines while minimizing false positives, and its recall demonstrates strong ability to capture all executed lines without missing any. Notably, half of the test dataset includes solutions from the CodeNet Project, published online since 2022, potentially benefiting LLMs if trained on this data. Despite this, CodeFlow still outperforms them.
IX-B3 Incomplete Code
To our knowledge, no runtime error dataset exists for incomplete Python code snippets paired with their complete versions (to build ground truth). To evaluate CodeFlow’s capability with incomplete code—specifically without built-in and external library imports—we trained and tested it on a dataset where all import statements and method/class declarations were removed. This preserved the code’s semantic integrity while retaining function/API calls. Despite the absence of imports, CodeFlow outperformed other models, as shown in Table I, due to its ability to learn library semantics during training and predict code behavior based on nodes using library functions. Similar performance was observed in LLMs (GPT-4o, Claude, and Gemini) due to their extensive pre-training on library semantics. CodeExecutor, which tracks intermediate execution values, and CFGNN, which models control flow, also performed well without import information.
@@ IX-B3 Incomplete Code
IX-B4 Continuity in Predicting Code Coverage
Continuity means a model should predict nodes or lines of code that are sequential or logically connected, rather than skipping intermediate nodes. This issue is common in existing methods. CodeExecutor, which relies on exact execution traces during training, often misses dependencies between lines. As shown in Fig. 6a, CodeExecutor skips a node and jumps directly to the next, creating a discontinuous CFG path.
@@ IX-B4 Continuity in Predicting Code Coverage
LLMs like GPT-4o also struggle with continuity, often failing to capture dynamic relations between lines and skipping critical steps, such as the elif in an if-elif-else structure (Fig. 6c). This happens because LLMs rely on next-token probabilities, predicting the most likely token based on prior context. As a result, they may misinterpret code structure, especially dynamic behaviors and state changes across iterations. Consequently, despite high precision, recall, and F1 scores, their exact-matching accuracy is lower.
@@ X-A Data Collection and Evaluation Metrics
We used a dataset in addition to that in RQ1. Specifically, we used the FixEval dataset, which comprises 2,066 unique problems with 277,262 submissions of Python code snippets. From this dataset, test cases were obtained for 800 problems from the CodeNet dataset [20]. Each of these snippets, when executed with its respective input, leads to a runtime error. After filtering, we obtained 6,437 submissions across the 800 problems. This combined dataset, referred to as the Complete Runtime-Error Dataset, includes both code snippets that terminate normally and those that encounter runtime errors.
In addition to the metrics in RQ1, we use Runtime Error Detection Accuracy (EDA) to measure the accuracy of a model correctly predicting if a snippet has an error.
@@ X-B Empirical Results
As seen in Table II, Claude achieved the highest precision of 0.98, indicating its strong ability to correctly identify runtime errors when they are present. High precision means Claude makes very few false positive predictions, thus showing its accuracy in pinpointing real runtime errors. However, Claude’s recall score of 0.89, while still respectable, is lower than that of GPT-4o. This suggests that Claude may miss some runtime errors, indicating that it is more conservative in error detection.
In contrast, GPT-4o achieved the highest recall of 0.99, showing its effectiveness in identifying nearly all runtime errors. GPT-4o’s precision score of 0.71 indicates a higher rate of false positives compared to Claude, meaning it sometimes incorrectly flags non-buggy code snippets as erroneous. This suggests that GPT-4o could have more false alarms.
@@
As seen, CodeFlow exhibits a balanced performance with high scores across all metrics: a precision of 0.96, a recall of 0.94, and an F1-score of 0.95. This balance indicates that CodeFlow not only accurately detects a high proportion of actual runtime errors but also minimizes false positives. The overall accuracy of 97.51% shows superior capability in statically identifying runtime errors without execution.